### v2

@@ 4.3 Alignment with Human Preferences
One of the most impactful applications of reinforcement learning (RL) in video generation is post-training alignment. This stage fine-tunes pretrained generators to better match subjective human preferences. Alignment methods vary in how they incorporate reward signals—ranging from explicit policy optimization to gradient-based preference modeling.
Policy Optimization: A core group of methods uses formal policy optimization to align generative models. While early RL algorithms struggled with stability, recent approaches offer more robust performance. Group Relative Policy Optimization (GRPO) is one such method. DanceGRPO [28] applies GRPO to a variety of visual generation tasks, including text-to-image, text-to-video, and image-to-video synthesis. It reformulates both diffusion sampling and rectified flows as stochastic differential equations (SDEs), allowing GRPO to operate across architectures and training paradigms. It supports multiple foundation models (e.g., Stable Diffusion, HunyuanVideo) and various reward types, such as aesthetic quality, alignment, motion consistency, and binary feedback. This flexibility enables GRPO to effectively instill structured knowledge during post-training.
@@
VideoDPO [140] is an early implementation that builds large-scale preference datasets using automated scoring heuristics. HuViDPO [143] extends this framework to text-to-video generation. It proposes a structured loss and architectural refinements—such as First-Frame Conditioning and Sparse Causal Attention—to improve visual quality and temporal consistency.
@@
To support fine-grained alignment, several methods decompose preferences into interpretable dimensions. VisionReward [142] uses Multi-Objective Preference Optimization (MPO) to balance temporal consistency, text fidelity, and realism. HALO [145] further introduces Granular-DPO, which applies patch-level feedback to correct localized artifacts. DenseDPO [152] improves structural alignment by denoising corrupted videos and using segment-level annotations for supervision. This reduces annotation cost while preserving quality.
@@
Advances in preference data sourcing have also driven DPO’s scalability. RDPO [7] uses physics-based heuristics to automatically generate preference pairs from real videos. For example, preferring a ball that falls over one that floats. This approach encodes physical plausibility without human labeling.
@@
Some studies improve the learning dynamics of DPO itself. OnlineVPO [141] proposes an online variant that continuously samples, ranks, and updates the model in real time. This reduces temporal artifacts and improves subject consistency without relying on static datasets.
@@
Moreover, DPO has also been integrated into hybrid training pipelines. SePPO [133] and EchoMimicV3 [153] alternate between supervised learning and preference optimization, combining the stability of supervised training with the alignment precision of DPO. VPO [154] extends this idea by optimizing user prompts, rather than generation models, using DPO. SkyReelsV2 [147] implements a multi-stage pipeline involving supervised finetuning, motion-specific DPO, infinite-length generation, and final polishing. This structured framework enables open-source cinematic video synthesis.
@@
New formulations of the DPO objective have also emerged. AAVG [146] introduces Gap-Aware Preference Optimization (GAPO), which weights training samples based on preference confidence. IDOI [138] combines DPO and Reward-Weighted Regression (RWR) in a unified probabilistic framework, using binary vision-language feedback to optimize dynamic object interactions. IPO [144] proposes Iterative Preference Optimization, an RL-based framework that alternates between evaluation and critic-guided updates. This improves semantic fidelity, motion realism, and aesthetic appeal. In contrast, Diffusion-NPO [149] focuses on undesirable outputs. It introduces Negative Preference Optimization (NPO), which trains models to suppress unwanted behaviors using inverted preference labels. The negatively aligned model is used as the unconditional branch in classifier-free guidance, enhancing separation between high- and low-quality generations. This strategy improves visual quality, robustness, and human alignment across tasks.
@@
InstructVideo [131] represents an early instance of this paradigm. It formulates video generation as an editing task and leverages pretrained image-level reward models to evaluate outputs. This lightweight strategy reduces computational overhead by iteratively applying reward feedback without full model retraining. VADER [132] expands this framework by integrating multiple expert models—such as CLIP for semantic alignment and YOLO for object detection—into a unified, differentiable reward function. This enables gradient-based optimization across multiple quality dimensions, improving output fidelity and relevance.
In contrast to these gradient-based methods, Free2Guide [137] adopts a gradient-free approach inspired by optimal control. It uses black-box reward signals—such as evaluations from vision-language models—to guide inference-time decisions. This plug-and-play strategy conditions generation on reward signals without modifying the generator’s parameters.
@@ 4.4 Physical Consistency and World Modeling
Phys-AR [148] is a pioneering effort in this direction. It integrates symbolic reasoning with reinforcement learning in a post-training setting. The model first converts video frames into symbolic tokens using a Diffusion Timestep Tokenizer (DDT). It then frames generation as a token-level MDP and applies GRPO to refine the generation policy. A key contribution of Phys-AR lies in its reward design, which encodes physical laws such as velocity consistency and mass-informed motion. This encourages the model to generate trajectories that exhibit uniform motion, parabolic arcs, and realistic collision responses.
RLVR-World [150] offers a complementary approach, addressing a fundamental limitation of MLE in video generation. MLE often produces outputs that are blurry or semantically unstable, as it does not optimize for perceptual quality. RLVR-World introduces a reinforcement learning-based fine-tuning scheme that directly targets perceptual rewards, including LPIPS, SSIM, and task accuracy. GRPO is again employed to align the generation policy with these metrics. Importantly, this framework extends beyond video, demonstrating applicability to language modeling and other generative tasks.
@@ 4.5 Inference-Time Reinforcement Learning
InfLVG [151] demonstrates this paradigm by incorporating GRPO directly into the inference loop. Instead of relying on static context selection, InfLVG formulates generation as a sequential decision process. At each step, the model samples a set of candidate continuations and scores them using a composite reward function. This reward balances three key aspects: (1) content consistency, measured by face identity similarity; (2) prompt relevance, assessed via CLIP-based text–video alignment; and (3) artifact suppression, penalized using a vision–language classifier. These scores are used to compute relative advantages among candidates. GRPO then updates the policy via a clipped objective, guiding a Plackett–Luce top-K sampling procedure to select the most appropriate context tokens. This inference-time optimization enables adaptive context selection while keeping the computational budget fixed. As a result, InfLVG achieves significant length extension—up to 9× longer—while improving visual sharpness and temporal consistency.
@@
VC2+VideoDPO [140] - 52.29 99.00 92.18 95.69 63.18
Turbo+VideoDPO [140] - 51.98 94.00 88.85 96.10 68.98
CogVid+VideoDPO [140] - 54.04 81.00 88.64 94.67 58.64
VisionReward [142] - 71.54 98.40 - - -
Turbo-v1+HALO [145] 72.07 54.97 95.00 - - -
Turbo-v2+HALO [145] 69.11 67.5 97.6 - - -
CogVidX-2B+HALO [145] 61.90 72.91 98.00 - - -
RDPO [7] 65.11 - - 99.27 97.04 55.01
CogVidX-2B+VPO [154] - 70.17 99.00 - - -
CogVidX-5B+VPO [154] - 75.73 99.60 - - -
AAVG [146] 68.98 - - 99.13 95.20 54.28
OnlineVPO [141] 67.36 - - 99.36 97.58 55.37
InstructVideo [131] 70.09 - - 96.76 96.45 50.01
VADER [132] 66.08 - - 98.89 95.53 53.43
CogVidX-2B+IPO [144] 62.87 - - 98.17 96.79 62.31
@@ 5 RL for 3D Content Generation
The integration of reinforcement learning (RL) into 3D content generation has introduced a powerful paradigm that blends sequential decision-making with geometric modeling. This synergy enables more adaptive, controllable, and preference-aligned synthesis of complex 3D structures. As illustrated in Figure 2, recent research efforts in this domain are broadly distributed across multiple subfields, including Text-to-NeRF/3D Gaussian Splatting, 3D diffusion modeling, multi-view consistency optimization, human motion generation, point cloud reconstruction, and other emerging directions. The balanced distribution among these topics underscores the breadth and growing maturity of RL applications in 3D generation. Unlike earlier stages dominated by a single task or representation, current trends suggest that RL is being explored as a general-purpose optimization framework across the entire 3D content generation pipeline.
@@
Point Cloud-Based 3D Generation: Akizuki et al. [155] were among the first to explore RL for 3D generation. Their method voxelizes 3D models to define spatial boundaries. The action space corresponds to voxel extension directions, while the reward encourages topological validity and penalizes self-intersections. This setup enables the agent to generate diverse, structurally sound 3D objects by learning motion trajectories within voxel grids.
Recent works extend this idea to mesh generation. DeepMesh [156] employs Direct Preference Optimization (DPO) in 3D autoregressive models to align generation with human preferences. However, this method depends on costly manual annotation of preference pairs. Mesh-RFT [157] proposes a topology-aware scoring scheme using automated metrics—Boundary Edge Ratio (BER) and Topology Score (TS)—to assess mesh quality at both object and face levels. It further introduces M-DPO and quality-aware masking to refine flawed regions and capture local geometric variations.
Depth Image-Based 3D Generation: Zhang et al.[158] propose a deep RL method based on A3C to reconstruct complete 3D scenes from RGB-D inputs, even under heavy occlusion. The model selects optimal viewpoints via a learned policy and applies volume-guided inpainting for progressive semantic completion. Similarly, Lin et al.[10] use depth images as structural references and train two RL agents—one for modifying primitives, the other for adjusting vertices. Their reward is based on the change in intersection-over-union (IoU), enabling precise geometric editing.
@@
DreamReward [9] represents RL-based frameworks and constructs a large-scale preference-labeled dataset of 25,304 prompt–3D asset pairs with human feedback scores. It trains a reward model, Reward3D, to evaluate generation quality. The sampling process is formulated as an RL task, where actions correspond to updates of the 3D representation, and the reward is derived from Reward3D. By integrating this learned reward with Score Distillation Sampling (SDS) [159], the method improves both alignment and fidelity of 3D outputs.
However, DreamReward’s reliance on a dedicated reward model introduces scalability limitations. DreamDPO [160] addresses this by applying Direct Preference Optimization (DPO) with large pretrained vision-language models acting as zero-shot reward functions. During training, image renderings of 3D assets are ranked using textual instructions, and these rankings guide preference-based updates to the 3D representation. This framework removes the need for manual reward supervision while leveraging the generalization ability of large multimodal models. By framing the process as offline preference optimization, DreamDPO aligns 3D generation with RL principles while avoiding the instability of dynamic exploration.
To mitigate the need for preference-labeled data, DreamCS [161] introduces a scalable alternative using unpaired 3D meshes and a Cauchy–Schwarz (CS) divergence-based reward objective. It trains a geometry-aware reward model, RewardCS, on this unpaired dataset, and applies reward-guided sampling to iteratively improve the 3D output. Although not explicitly framed as an RL agent-environment loop, DreamCS still adheres to the RL paradigm by defining a reward-driven optimization landscape with structured feedback.
Recent methods have also integrated RL into the training of large-scale 3D diffusion models. GPLD3D [162] formulates two domain-specific reward functions—geometric feasibility and physical stability—and incorporates them into the generation loop to guide high-quality asset synthesis. In a more generalizable formulation, Nabla-R2D3 [163] proposes a novel RL alignment framework that bridges 2D and 3D reward spaces. It uses a pretrained 2D reward model to guide 3D generation via a probabilistic refinement mechanism, Nabla-GFlowNet, which transforms 2D signals into structured 3D rewards. This approach highlights a key strength of RL in 3D generation: the ability to optimize across modalities using proxy signals, even in the absence of explicit 3D supervision.
@@ 5.2 Multi-View Model Enhancement
Carve3D [164] introduces RL Fine-Tuning (RLFT) into multi-view diffusion models. Starting from a pretrained text-to-image model, Carve3D applies RLFT using a novel Multi-view Reconstruction Consistency (MRC) metric. This metric compares images generated from multiple viewpoints with views rendered from a NeRF reconstructed by a large sparse-view model (LRM). The MRC score serves as the reward signal, guiding the diffusion model to produce outputs that are both visually plausible and geometrically consistent across views.
HFDream [165] advances this direction by constructing a large-scale dataset of multi-view image–text pairs annotated with human feedback. A reward model is trained to assess the viewpoint accuracy of generated images relative to textual prompts. This model is then used to fine-tune a diffusion generator via reward-weighted loss, directly aligning multi-view outputs with both human preferences and geometric fidelity in text-to-3D scenarios.
Pushing RL integration further into the image-to-3D domain, MVReward [166] collects a dataset of multi-view RGB images, normal maps, and associated human preference annotations. It introduces a multi-view preference learning strategy that combines reward loss—derived from a learned preference model—with standard pretraining objectives. This hybrid RL formulation improves the structural realism and viewpoint alignment of reconstructed 3D shapes.
To alleviate the tension between high computational cost and generation quality in few-step text-to-multi-view models, Zhang et al. [167] propose a novel RL fine-tuning framework that reformulates the task as a constrained single-view policy optimization problem. Their method introduces ZMV-Sampling, which enhances both cross-view and text-image alignment, and MV-ZigAL, a reward-advantage learning scheme. MV-ZigAL uses advantage-weighted policy updates to internalize reward-driven improvements, enabling the base sampling policy to generate coherent multi-view images with fewer inference steps.

### v3

@@ 4.3 Alignment with Human Preferences
One of the most impactful applications of reinforcement learning (RL) in video generation is post-training alignment. This stage fine-tunes pretrained generators to better match subjective human preferences. Alignment methods vary in how they incorporate reward signals—ranging from explicit policy optimization to gradient-based preference modeling. The design and quality of these reward signals are therefore fundamental. Table 3 provides an overview of representative reward models commonly used to guide such alignment processes.
Policy Optimization: A core group of methods uses formal policy optimization to align generative models. While early RL algorithms struggled with stability, recent approaches offer more robust performance. Group Relative Policy Optimization (GRPO) is one such method. DanceGRPO [dancegrpo] applies GRPO to a variety of visual generation tasks, including text-to-image, text-to-video, and image-to-video synthesis. It reformulates both diffusion sampling and rectified flows as stochastic differential equations (SDEs), allowing GRPO to operate across architectures and training paradigms. It supports multiple foundation models (e.g., Stable Diffusion, HunyuanVideo) and various reward types, such as aesthetic quality, alignment, motion consistency, and binary feedback. This flexibility enables GRPO to effectively instill structured knowledge during post-training.
@@
VideoDPO [videodpo] is an early implementation that builds large-scale preference datasets using automated scoring heuristics. HuViDPO [jiang2025huvidpo] extends this framework to text-to-video generation. It proposes a structured loss and architectural refinements—such as First-Frame Conditioning and Sparse Causal Attention—to improve visual quality and temporal consistency.
@@
To support fine-grained alignment, several methods decompose preferences into interpretable dimensions. VisionReward [visionreward] uses Multi-Objective Preference Optimization (MPO) to balance temporal consistency, text fidelity, and realism. HALO [wang2025harness] further introduces Granular-DPO, which applies patch-level feedback to correct localized artifacts. DenseDPO [densedpo] improves structural alignment by denoising corrupted videos and using segment-level annotations for supervision. This reduces annotation cost while preserving quality.
@@
Advances in preference data sourcing have also driven DPO’s scalability. RDPO [rdpo] uses physics-based heuristics to automatically generate preference pairs from real videos. For example, preferring a ball that falls over one that floats. This approach encodes physical plausibility without human labeling.
@@
Some studies improve the learning dynamics of DPO itself. OnlineVPO [zhang2024onlinevpo] proposes an online variant that continuously samples, ranks, and updates the model in real time. This reduces temporal artifacts and improves subject consistency without relying on static datasets.
@@
Moreover, DPO has also been integrated into hybrid training pipelines. SePPO [sspo] and EchoMimicV3 [echomimicv3] alternate between supervised learning and preference optimization, combining the stability of supervised training with the alignment precision of DPO. VPO [vpo] extends this idea by optimizing user prompts, rather than generation models, using DPO. SkyReelsV2 [skyreels] implements a multi-stage pipeline involving supervised finetuning, motion-specific DPO, infinite-length generation, and final polishing. This structured framework enables open-source cinematic video synthesis.
@@
New formulations of the DPO objective have also emerged. AAVG [zhu2025aligning] introduces Gap-Aware Preference Optimization (GAPO), which weights training samples based on preference confidence. IDOI [furuta2024improving] combines DPO and Reward-Weighted Regression (RWR) in a unified probabilistic framework, using binary vision-language feedback to optimize dynamic object interactions. IPO [yang2025ipo] proposes Iterative Preference Optimization, an RL-based framework that alternates between evaluation and critic-guided updates. This improves semantic fidelity, motion realism, and aesthetic appeal. In contrast, Diffusion-NPO [wang2025diffusion] focuses on undesirable outputs. It introduces Negative Preference Optimization (NPO), which trains models to suppress unwanted behaviors using inverted preference labels. The negatively aligned model is used as the unconditional branch in classifier-free guidance, enhancing separation between high- and low-quality generations. This strategy improves visual quality, robustness, and human alignment across tasks.
@@
InstructVideo [instructvideo] represents an early instance of this paradigm. It formulates video generation as an editing task and leverages pretrained image-level reward models to evaluate outputs. This lightweight strategy reduces computational overhead by iteratively applying reward feedback without full model retraining. VADER [vader] expands this framework by integrating multiple expert models—such as CLIP for semantic alignment and YOLO for object detection—into a unified, differentiable reward function. This enables gradient-based optimization across multiple quality dimensions, improving output fidelity and relevance.
In contrast to these gradient-based methods, Free2Guide [kim2024free] adopts a gradient-free approach inspired by optimal control. It uses black-box reward signals—such as evaluations from vision-language models—to guide inference-time decisions. This plug-and-play strategy conditions generation on reward signals without modifying the generator’s parameters.
@@ 4.4 Physical Consistency and World Modeling
Phys-AR [physar] is a pioneering effort in this direction. It integrates symbolic reasoning with reinforcement learning in a post-training setting. The model first converts video frames into symbolic tokens using a Diffusion Timestep Tokenizer (DDT). It then frames generation as a token-level MDP and applies GRPO to refine the generation policy. A key contribution of Phys-AR lies in its reward design, which encodes physical laws such as velocity consistency and mass-informed motion. This encourages the model to generate trajectories that exhibit uniform motion, parabolic arcs, and realistic collision responses.
RLVR-World [rlvrworld] offers a complementary approach, addressing a fundamental limitation of MLE in video generation. MLE often produces outputs that are blurry or semantically unstable, as it does not optimize for perceptual quality. RLVR-World introduces a reinforcement learning-based fine-tuning scheme that directly targets perceptual rewards, including LPIPS, SSIM, and task accuracy. GRPO is again employed to align the generation policy with these metrics. Importantly, this framework extends beyond video, demonstrating applicability to language modeling and other generative tasks.
@@ 4.5 Inference-Time Reinforcement Learning
InfLVG [inflvg] demonstrates this paradigm by incorporating GRPO directly into the inference loop. Instead of relying on static context selection, InfLVG formulates generation as a sequential decision process. At each step, the model samples a set of candidate continuations and scores them using a composite reward function. This reward balances three key aspects: (1) content consistency, measured by face identity similarity; (2) prompt relevance, assessed via CLIP-based text–video alignment; and (3) artifact suppression, penalized using a vision–language classifier. These scores are used to compute relative advantages among candidates. GRPO then updates the policy via a clipped objective, guiding a Plackett–Luce top-K sampling procedure to select the most appropriate context tokens. This inference-time optimization enables adaptive context selection while keeping the computational budget fixed. As a result, InfLVG achieves significant length extension—up to 9× longer—while improving visual sharpness and temporal consistency.
In summary, we provide a comparative analysis of representative approaches. Owing to the heterogeneity of experimental settings and the variability of reported evaluation protocols, we restrict our discussion to methods that have been quantitatively assessed on the VBench benchmark. The corresponding evaluation results are reported in Table 4.
@@
VC2+VideoDPO [videodpo] - 52.29 99.00 92.18 95.69 63.18
Turbo+VideoDPO [videodpo] - 51.98 94.00 88.85 96.10 68.98
CogVid+VideoDPO [videodpo] - 54.04 81.00 88.64 94.67 58.64
VisionReward [visionreward] - 71.54 98.40 - - -
Turbo-v1+HALO [wang2025harness] 72.07 54.97 95.00 - - -
Turbo-v2+HALO [wang2025harness] 69.11 67.5 97.6 - - -
CogVidX-2B+HALO [wang2025harness] 61.90 72.91 98.00 - - -
RDPO [rdpo] 65.11 - - 99.27 97.04 55.01
CogVidX-2B+VPO [vpo] - 70.17 99.00 - - -
CogVidX-5B+VPO [vpo] - 75.73 99.60 - - -
AAVG [zhu2025aligning] 68.98 - - 99.13 95.20 54.28
OnlineVPO [zhang2024onlinevpo] 67.36 - - 99.36 97.58 55.37
InstructVideo [instructvideo] 70.09 - - 96.76 96.45 50.01
VADER [vader] 66.08 - - 98.89 95.53 53.43
CogVidX-2B+IPO [yang2025ipo] 62.87 - - 98.17 96.79 62.31
@@ 5 RL for 3D Content Generation
The integration of reinforcement learning (RL) into 3D content generation has introduced a powerful paradigm that blends sequential decision-making with geometric modeling. This synergy enables more adaptive, controllable, and preference-aligned synthesis of complex 3D structures. As illustrated in Figure 3, recent research efforts in this domain are broadly distributed across multiple subfields, including Text-to-NeRF/3D Gaussian Splatting, 3D diffusion modeling, multi-view consistency optimization, human motion generation, point cloud reconstruction, and other emerging directions. The balanced distribution among these topics underscores the breadth and growing maturity of RL applications in 3D generation. Unlike earlier stages dominated by a single task or representation, current trends suggest that RL is being explored as a general-purpose optimization framework across the entire 3D content generation pipeline. An overview of recent representative studies is summarized in Table LABEL:tab:3d_table.
@@
Point Cloud-Based 3D Generation: Akizuki et al. [akizuki2020generative] were among the first to explore RL for 3D generation. Their method voxelizes 3D models to define spatial boundaries. The action space corresponds to voxel extension directions, while the reward encourages topological validity and penalizes self-intersections. This setup enables the agent to generate diverse, structurally sound 3D objects by learning motion trajectories within voxel grids.
Recent works extend this idea to mesh generation. DeepMesh [zhao2025deepmesh] employs Direct Preference Optimization (DPO) in 3D autoregressive models to align generation with human preferences. However, this method depends on costly manual annotation of preference pairs. Mesh-RFT [liu2025mesh] proposes a topology-aware scoring scheme using automated metrics—Boundary Edge Ratio (BER) and Topology Score (TS)—to assess mesh quality at both object and face levels. It further introduces M-DPO and quality-aware masking to refine flawed regions and capture local geometric variations.
Depth Image-Based 3D Generation: Zhang et al.[zhang2023point] propose a deep RL method based on A3C to reconstruct complete 3D scenes from RGB-D inputs, even under heavy occlusion. The model selects optimal viewpoints via a learned policy and applies volume-guided inpainting for progressive semantic completion. Similarly, Lin et al.[lin2020modeling] use depth images as structural references and train two RL agents—one for modifying primitives, the other for adjusting vertices. Their reward is based on the change in intersection-over-union (IoU), enabling precise geometric editing.
DreamReward [ye2024dreamreward] represents RL-based frameworks and constructs a large-scale preference-labeled dataset of 25,304 prompt–3D asset pairs with human feedback scores. It trains a reward model, Reward3D, to evaluate generation quality. The sampling process is formulated as an RL task, where actions correspond to updates of the 3D representation, and the reward is derived from Reward3D. By integrating this learned reward with Score Distillation Sampling (SDS) [poole2022dreamfusion], the method improves both alignment and fidelity of 3D outputs.
However, DreamReward’s reliance on a dedicated reward model introduces scalability limitations. DreamDPO [zhou2025dreamdpo] addresses this by applying Direct Preference Optimization (DPO) with large pretrained vision-language models acting as zero-shot reward functions. During training, image renderings of 3D assets are ranked using textual instructions, and these rankings guide preference-based updates to the 3D representation. This framework removes the need for manual reward supervision while leveraging the generalization ability of large multimodal models. By framing the process as offline preference optimization, DreamDPO aligns 3D generation with RL principles while avoiding the instability of dynamic exploration.
To mitigate the need for preference-labeled data, DreamCS [zou2025dreamcs] introduces a scalable alternative using unpaired 3D meshes and a Cauchy–Schwarz (CS) divergence-based reward objective. It trains a geometry-aware reward model, RewardCS, on this unpaired dataset, and applies reward-guided sampling to iteratively improve the 3D output. Although not explicitly framed as an RL agent-environment loop, DreamCS still adheres to the RL paradigm by defining a reward-driven optimization landscape with structured feedback.
Recent methods have also integrated RL into the training of large-scale 3D diffusion models. GPLD3D [dong2024gpld3d] formulates two domain-specific reward functions—geometric feasibility and physical stability—and incorporates them into the generation loop to guide high-quality asset synthesis. In a more generalizable formulation, Nabla-R2D3 [liu2025nabla] proposes a novel RL alignment framework that bridges 2D and 3D reward spaces. It uses a pretrained 2D reward model to guide 3D generation via a probabilistic refinement mechanism, Nabla-GFlowNet, which transforms 2D signals into structured 3D rewards. This approach highlights a key strength of RL in 3D generation: the ability to optimize across modalities using proxy signals, even in the absence of explicit 3D supervision.
@@ 5.2 Multi-View Model Enhancement
Carve3D [xie2024carve3d] introduces RL Fine-Tuning (RLFT) into multi-view diffusion models. Starting from a pretrained text-to-image model, Carve3D applies RLFT using a novel Multi-view Reconstruction Consistency (MRC) metric. This metric compares images generated from multiple viewpoints with views rendered from a NeRF reconstructed by a large sparse-view model (LRM). The MRC score serves as the reward signal, guiding the diffusion model to produce outputs that are both visually plausible and geometrically consistent across views.
HFDream [choi2024hfdream] advances this direction by constructing a large-scale dataset of multi-view image–text pairs annotated with human feedback. A reward model is trained to assess the viewpoint accuracy of generated images relative to textual prompts. This model is then used to fine-tune a diffusion generator via reward-weighted loss, directly aligning multi-view outputs with both human preferences and geometric fidelity in text-to-3D scenarios.
Pushing RL integration further into the image-to-3D domain, MVReward [wang2025mvreward] collects a dataset of multi-view RGB images, normal maps, and associated human preference annotations. It introduces a multi-view preference learning strategy that combines reward loss—derived from a learned preference model—with standard pretraining objectives. This hybrid RL formulation improves the structural realism and viewpoint alignment of reconstructed 3D shapes.
To alleviate the tension between high computational cost and generation quality in few-step text-to-multi-view models, Zhang et al. [zhang2025refining] propose a novel RL fine-tuning framework that reformulates the task as a constrained single-view policy optimization problem. Their method introduces ZMV-Sampling, which enhances both cross-view and text-image alignment, and MV-ZigAL, a reward-advantage learning scheme. MV-ZigAL uses advantage-weighted policy updates to internalize reward-driven improvements, enabling the base sampling policy to generate coherent multi-view images with fewer inference steps.